15. A body projected upward from the level ground at certain angle, if the total time taken is 6.3sec with an initial velocity of 40m/s. Find (i) (ii) (iii) (iv) The angle of projection How far from the starting point will it strike the ground How height will it go  Max. Range  


1
Expert's answer
2021-06-11T11:21:28-0400

1)

"\alpha=\arcsin \frac{gt}{2v}=50.5°,"

2)

"l=\frac{v^2\sin 2\alpha}{g}=153~m,"

3)

"h=\frac{v^2\sin^2\alpha}{2g}=49~m,"

4)

"s=l=153~m."